# Quantum One-Time Programs, Revisited

### v1

@@ D.3 Knowledge of Preimage for Expanding Random Oracles
Consider the adversary $B$ which attempts to find $k$ input/output pairs by running $A$ to obtain $\mathbf{y}$, then guessing a uniform $\mathbf{x}\in\mathcal{X}^{k}$ and outputting $(\mathbf{x},\mathbf{y})$. Denote $p_{B}$ as the probability that it outputs $(\mathbf{x},\mathbf{y})$ such that $G(x_{i})=y_{i}$ for all $i$. Since $\mathbf{x}$ is independent of $A$, we have $p_{B}\geq p/|\mathcal{X}|^{k}$.
Now consider running $B$ with a compressed oracle, then measuring the compressed oracle to obtain a database $D$. Note that this is the same distribution over databases as running $A$. Denote $p^{\prime}_{B}$ as the probability that it successfully outputs $(\mathbf{x},\mathbf{y})$ such that $(x_{i},y_{i})\in D$. We may decompose the corresponding event into two mutually exclusive components:
  • Let $E^{\prime}_{B,+}$ be the event that $(x_{i},y_{i})\in D$ for all $i$ and $D$ contains a collision $(x^{*}_{0},y^{*})$ and $(x^{*}_{1},y^{*})$.
  • Let $E^{\prime}_{B,1}$ be the event that $(x_{i},y_{i})\in D$ for all $i$ and $D$ does not contain a collision.
By definition, $p^{\prime}_{B}=\Pr[E^{\prime}_{B,+}]+\Pr[E^{\prime}_{B,1}]$. We may define analogous events $E^{\prime}_{A,+}$ and $E^{\prime}_{A,1}$ when $A$ is run and the database is measured, where the first condition is changed to the existence of $x_{i}$, rather than requiring $A$ to find it. $\Pr[E^{\prime}_{B,+}]$ and $\Pr[E^{\prime}_{A,+}]$ are both bounded by the probability of $D$ containing a collision, which Lemma 3.5 bounds by $O(q^{3}/|\mathcal{Y}|)$. Furthermore, since there is a unique “solution” to $\mathbf{y}$ in the event $E^{\prime}_{A,1}$, we have $\Pr[E^{\prime}_{B,1}]=\Pr[E^{\prime}_{A,1}]/|\mathcal{X}|^{k}$. Thus we can relate $p^{\prime}_{A}$ and $p^{\prime}_{B}$ by $p^{\prime}_{B}\leq O(q^{3}/|\mathcal{Y}|)+\Pr[E^{\prime}_{A,1}]/|\mathcal{X}|^{k}\leq O(q^{3}/|\mathcal{Y}|)+p^{\prime}_{A}/|\mathcal{X}|^{k}$.

### v2

@@ C.3 Knowledge of Preimage for Expanding Random Oracles
Consider the adversary $B$ which attempts to find $k$ input/output pairs by running $A$ to obtain $\mathbf{x^{(1)}}$ and $\mathbf{y}$, then guessing a uniform $\mathbf{x^{(2)}}\in\mathcal{X}_{2}^{k}$ to construct the vector $\mathbf{x}=\left(x_{1}^{(1)}\|x_{1}^{(2)},\dots,x_{k}^{(1)}\|x_{k}^{(2)}\right)$ and outputting $(\mathbf{x},\mathbf{y})$. Denote $p_{B}$ as the probability that it outputs $(\mathbf{x},\mathbf{y})$ such that $G(x_{i}^{(1)},x_{i}^{(2)})=y_{i}$ for all $i$. Since $\mathbf{x^{(2)}}$ is independent of $A$, we have $p_{B}\geq p/|\mathcal{X}_{2}|^{k}$.
Now consider running $B$ with a compressed oracle, then measuring the compressed oracle to obtain a database $D$. Note that this is the same distribution over databases as running $A$. Denote $p^{\prime}_{B}$ as the probability that it successfully outputs $(\mathbf{x},\mathbf{y})$ such that $(x_{i}^{(1)}\|x_{i}^{(2)},y_{i})\in D$. We may decompose the corresponding event into two mutually exclusive components:
  • Let $E^{\prime}_{B,+}$ be the event that $(x_{i}^{(1)}\|x_{i}^{(2)},y_{i})\in D$ for all $i$ and $D$ contains a collision $(x^{*}_{0},y^{*})$ and $(x^{*}_{1},y^{*})$.
  • Let $E^{\prime}_{B,1}$ be the event that $(x_{i}^{(1)}\|x_{i}^{(2)},y_{i})\in D$ for all $i$ and $D$ does not contain a collision.
By definition, $p^{\prime}_{B}=\Pr[E^{\prime}_{B,+}]+\Pr[E^{\prime}_{B,1}]$. We may define analogous events $E^{\prime}_{A,+}$ and $E^{\prime}_{A,1}$ when $A$ is run and the database is measured, where the first condition is changed to the existence of $x_{i}$, rather than requiring $A$ to find it. $\Pr[E^{\prime}_{B,+}]$ and $\Pr[E^{\prime}_{A,+}]$ are both bounded by the probability of $D$ containing a collision, which Lemma 3.5 bounds by $O(q^{3}/|\mathcal{Y}|)$. Furthermore, since there is a unique “solution” to $\mathbf{y}$ in the event $E^{\prime}_{A,1}$, we have $\Pr[E^{\prime}_{B,1}]=\Pr[E^{\prime}_{A,1}]/|\mathcal{X}_{2}|^{k}$. Thus we can relate $p^{\prime}_{A}$ and $p^{\prime}_{B}$ by $p^{\prime}_{B}\leq O(q^{3}/|\mathcal{Y}|)+\Pr[E^{\prime}_{A,1}]/|\mathcal{X}_{2}|^{k}\leq O(q^{3}/|\mathcal{Y}|)+p^{\prime}_{A}/|\mathcal{X}_{2}|^{k}$.
@@
Appendix D Additional Prelims
We give some additional preliminaries in this section.
D.1 NIZK
A NIZK for NP scheme should satisfy the following properties:
Correctness
A NIZK proof $(\mathsf{Setup},\mathsf{Delegate},\mathsf{Prove},\mathsf{Verify})$ is correct if there exists a negligible function $\mathsf{negl}(\cdot)$ such that for all $\lambda\in\mathbb{N}$, all $x\in L$, and all $w\in\mathcal{R}_{L}(x)$ it holds that
$\Pr[\mathsf{Verify}(\mathsf{CRS},\mathsf{Prove}(\mathsf{CRS},\mathsf{token},\ w,x),x)=1]=1-\mathsf{negl}(\lambda)$
where $(\mathsf{CRS},\mathsf{msk})\leftarrow\mathsf{Setup}(1^{\lambda}),\mathsf{token}\leftarrow\mathsf{Delegate}(\mathsf{msk})$.
Computational Soundness
A one-time NIZK proof $(\mathsf{Setup},\mathsf{Prove},\mathsf{Verify})$ is computationally sound if there exist a negligible function $\mathsf{negl}(\cdot)$ such that for all unbounded adversaries $\mathcal{A}$ and all $x\notin L$, it holds that:
$\Pr[\mathsf{Verify}(\mathsf{CRS},\pi\leftarrow\mathcal{A}(\mathsf{CRS},x,\mathsf{token}),x)=1]=\mathsf{negl}(\lambda)$
where $(\mathsf{CRS},\mathsf{msk})\leftarrow\mathsf{Setup}(1^{\lambda}),\mathsf{token}\leftarrow\mathsf{Delegate}(\mathsf{msk})$.
Computational Zero Knowledge
A one-time NIZK proof $(\mathsf{Setup},\mathsf{Delegate},\mathsf{Prove},\mathsf{Verify})$ is computationally zero-knowledge if there exists a simulator $S$ such that for all non-uniform QPT adversaries with quantum advice $\{\rho_{\lambda}\}_{\lambda\in\mathbb{N}}$, all statements $x\in L$ and all witnesses $w\in\mathcal{R}_{L}(x)$, it holds that $S(1^{\lambda},x)\approx_{c}\mathsf{Prove}(\mathsf{CRS},\mathsf{token},w,x)$ where $\mathsf{CRS}\leftarrow\mathsf{Setup}(1^{\lambda}),\mathsf{token}\leftarrow\mathsf{Delegate}(\mathsf{msk})$.